# Quantum Aspects of Natural Top Quark Condensation

### v2

@@ III.2 Quartic Coupling
$\approx-\frac{{g}^{4}_{0}\hat{\lambda}}{2M_{0}^{8}}\biggl{(}\!\int_{xy}\!\!({B}^{\dagger}(x,y){B}(x,y))^{2}\delta^{4}(x-y)\biggr{)}$ (70)
@@ III.3 Full Renormalized Action
The resulting action with the principal quantum effects considered here is then:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S^{\prime}=W(\Omega)+\!\int_{x}\biggl{(}[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]+[\overline{\psi}_{R}\;iD\!\!\!/\!\,\,\psi_{R}]\biggr{)}+\int_{xy}\;\biggl{(}-[B^{i}{}^{\dagger}D_{F}B_{i}]+g_{0}(B^{i\dagger}D_{F}C_{i}+h.c.)\biggr{)}$ (72)
where we recall eq.(29), $C_{i}(x,y)=\sqrt{N_{c}}M_{0}^{2}H_{i}(x,y)+[\overline{\psi}_{R}(y)\psi_{iL}(x)]_{f}$, and we have from the loop calculations:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!W(\Omega)=\!\int_{xyx^{\prime}y^{\prime}}\!\!\!\![B^{\dagger}(x,y){\cal{F}}(x,y,x^{\prime},y^{\prime})B(x^{\prime},y^{\prime})],+\!\int_{x...z^{\prime}}\!\!\!\![B^{\dagger}(x,y)B(x^{\prime},y^{\prime})]{\cal{G}}(x,...,z^{\prime})[B^{\dagger}(w,z)B(w^{\prime},z^{\prime})]+$...,
$=\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\int_{xy}\;B^{\dagger}(x,y)D_{F}(x-y)B(x,y)-\frac{g_{0}^{4}\hat{\lambda}}{2M_{0}^{8}}\!\int_{xy}\!\!({B}^{\dagger}(x,y){B}(x,y))^{2}\delta^{4}(x-y)$ (73)
Hence $S^{\prime}$ can be can be rewritten as:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S^{\prime}=\!\int_{x}\biggl{(}[\overline{\psi}_{R}iD\!\!\!/\!\,\,\psi_{R}]_{f}+[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]_{f}\biggr{)}+\!\int_{xy}\!\biggl{(}g_{0}B^{i\dagger}(x,y)D_{F}(x-y)C_{i}+h.c.\biggr{)}$
$-\!\int_{xy}\![B^{\dagger}(x,y)D_{F}(x-y)B(x,y)]\biggl{(}1-\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\biggr{)}-\frac{g_{0}^{4}\hat{\lambda}}{2M_{0}^{8}}\int_{xy}\![B^{\dagger}(x,y)B(x,y)]^{2}\delta^{4}(x-y)$ (74)
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S^{\prime}=\!\int_{x}\biggl{(}[\overline{\psi}_{R}iD\!\!\!/\!\,\,\psi_{R}]_{f}+[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]_{f}\biggr{)}+\!\int_{xy}\!\biggl{(}\overline{g}_{0}B^{\prime}_{i}{}^{\dagger}(x,y)D_{F}(x-y)C^{i}(x,y)+h.c.\biggr{)}$
@@
$W(\Psi,H)=\!\int_{x}\!\biggl{(}[\overline{\psi}_{L}(x)iD\!\!\!/\!\,\,\psi_{L}(x)]+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]\biggr{)}$
@@
$\overline{g}_{0}^{2}={g_{0}^{2}}\left(1-\frac{N_{c}g_{0}^{2}}{8\pi^{2}}\right)^{-1}$ (80)
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\overline{g}_{0}^{2}\!\int_{xy}\!\!\!\!\sqrt{N_{c}}M_{0}^{2}(H^{i\dagger}(x,y)D_{F}(x-y)[\overline{\psi}_{R}(y)\psi_{iL}(x)]_{f}+h.c.)\approx\overline{g}_{Y}\int d^{4}X\biggl{(}H^{i\dagger}(X)[\overline{\psi}_{R}(X)\psi_{iL}(X)]_{f}+h.c.\biggr{)}$ (81)
@@
$\overline{g}_{Y}=\overline{g}_{0}^{2}\sqrt{2N_{c}/J}\;\phi(0)$ (82)
@@
$W(\Psi,H)=\!\int_{x}\!\biggl{(}[\overline{\psi}_{L}(x)iD\!\!\!/\!\,\,\psi_{L}(x)]_{f}+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]_{f}\biggr{)}$
@@
$W(\Psi,H)=\!\int_{x}\!\biggl{(}[\overline{\psi}_{L}(x)iD\!\!\!/\!\,\,\psi_{L}(x)]_{f}+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]_{f}\biggr{)}$
@@
$\sim\widetilde{Z}\partial_{X}B^{\prime i\dagger}\partial_{X}B^{\prime}_{i}-M_{0}^{2}B^{\prime i\dagger}B^{\prime}_{i}\qquad\text{where,}\qquad\widetilde{Z}=\frac{\overline{g}_{0}^{2}N_{c}}{8\pi^{2}}\ln\left(\frac{M_{0}}{m}\right)$ (86)
@@ IV.1 Physical Consequences of the Improved Yukawa Coupling
$\frac{\overline{g}_{0c}^{2}N_{c}}{8\pi^{2}}=1.06940.\qquad\text{cf. NJL model:}\;\;\frac{\overline{g}_{0c}^{2}N_{c}}{8\pi^{2}}=1$ (88)
@@
$\phi(r)\approx\frac{ce^{-|\mu|r}}{r};\;\;\;\;\;\makebox{then we find,}\;\;\;\;\;\phi(0)=\frac{2\sqrt{2}}{\pi^{3/2}}\sqrt{\frac{|\mu|}{M_{0}}}=0.50795\sqrt{\frac{|\mu|}{M_{0}}}$ (89)
@@
we then predict $M_{0}=6.745\;\makebox{TeV}$. [^6] This was computed in main where we quoted $5.9$ TeV, which resulted from using the NJL critical value $N_{c}g_{0c}^{2}/8\pi^{2}=1$ rather than the true $N_{c}g_{0c}^{2}/8\pi^{2}=1.0694$. However, we think these are within the uncertainty of the leading order calculations for the SKG solutions, and does not take RG running effects, e.g., logarithms, into account. Likewise, using an approximate numerical solution to the SKG equation yields $M_{0}=5.23\;\makebox{TeV.}$, (this is sensitive to a fitting function, and likely underestimates $M_{0}$). Conservatively we would estimate $M_{0}$ to lie in a range of order $5<M_{0}<7$ TeV.
@@
$\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\approx\frac{1.0694}{1+1.0694}=0.51677$ (91)
@@ IV.2 Physical Consequences of the Improved Quartic Coupling
$\lambda\approx(\overline{g}_{Y}^{4})\frac{N_{c}}{4\pi^{2}}\ln\left(\frac{M_{0}}{|\mu|}\right)\approx 0.321\;\;\;\makebox{(cf., $0.25$ experiment.)};$ (93)
@@
$\lambda\approx(\overline{g}_{Y}^{4}-\overline{g}_{Y}^{2}\lambda-\lambda^{2})\frac{N_{c}}{4\pi^{2}}\ln\left(\frac{M_{0}}{|\mu|}\right)\approx 0.230\;\;\;\makebox{(cf., $0.25$ experiment.)};$ (94)
@@
$\lambda\approx(\overline{g}_{Y}^{4}-\overline{g}_{Y}^{2}\lambda)\frac{N_{c}}{4\pi^{2}}\ln\left(\frac{M_{0}}{\mu}\right)\approx 0.243\;\;\;\makebox{(cf., $0.25$ experiment.)};$ (95)
Note that since we only know $M_{0}$ to precision of order $\sim 6\;\pm\;1$ TeV, we expect the $\lambda$ prediction is uncertain by (at least) $\sim\pm 0.01$.
@@ Acknowledgments
I thank Bill Bardeen, Bogdan Dobrescu, Andreas Kronfeld and Julius Kuti for discussions and comments during the course of this work.

### v3

@@ III.2 Quartic Coupling
$\qquad\approx-\frac{{g}^{4}_{0}\hat{\lambda}}{2M_{0}^{8}}\biggl{(}\!\int_{xy}\!\!({B}^{\dagger}(x,y){B}(x,y))^{2}\delta^{4}(x-y)\biggr{)},$ (78)
@@ III.3 Full Renormalized Action
The resulting action with eq.(44) and the principal quantum effects considered here is then :
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S=S_{KTH}+W(\Omega)+\!\int_{x}\biggl{(}[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]_{f}+[\overline{\psi}_{R}\;iD\!\!\!/\!\,\,\psi_{R}]_{f}\biggr{)}+\int_{xy}\;\biggl{(}-[B^{i}{}^{\dagger}D_{F}B_{i}]+g_{0}(B^{i\dagger}D_{F}C_{i}+h.c.)\biggr{)}.$ (80)
We recall eq.(30), $C_{i}(x,y)=\sqrt{N_{c}}M_{0}^{2}H_{i}(x,y)+[\overline{\psi}_{R}(y)\psi_{iL}(x)]_{f}$, and we have from the loop calculations:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!W(\Omega)=\!\int_{xyx^{\prime}y^{\prime}}\!\!\!\![B^{\dagger}(x,y){\cal{F}}(x,y,x^{\prime},y^{\prime})B(x^{\prime},y^{\prime})]+\!\int_{x...z^{\prime}}\!\!\!\![B^{\dagger}(x,y)B(x^{\prime},y^{\prime})]{\cal{G}}(x,...,z^{\prime})[B^{\dagger}(w,z)B(w^{\prime},z^{\prime})]+$...,
$=\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\int_{xy}\;B^{\dagger}(x,y)D_{F}(x-y)B(x,y)-\frac{g_{0}^{4}\hat{\lambda}}{2M_{0}^{8}}\!\int_{xy}\!\!({B}^{\dagger}(x,y){B}(x,y))^{2}\delta^{4}(x-y).$ (81)
Hence $S$ can be can be rewritten as:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S=S_{KTH}+\!\int_{x}\biggl{(}[\overline{\psi}_{R}iD\!\!\!/\!\,\,\psi_{R}]_{f}+[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]_{f}\biggr{)}+\!\int_{xy}\!\biggl{(}g_{0}B^{i\dagger}(x,y)D_{F}(x-y)C_{i}+h.c.\biggr{)}$
$-\!\int_{xy}\![B^{\dagger}(x,y)D_{F}(x-y)B(x,y)]\biggl{(}1-\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\biggr{)}-\frac{g_{0}^{4}\hat{\lambda}}{2M_{0}^{8}}\int_{xy}\![B^{\dagger}(x,y)B(x,y)]^{2}\delta^{4}(x-y).$ (82)
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S=S_{KTH}+\!\int_{x}\biggl{(}[\overline{\psi}_{R}iD\!\!\!/\!\,\,\psi_{R}]_{f}+[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]_{f}\biggr{)}+\!\int_{xy}\!\biggl{(}\overline{g}_{0}B^{\prime}_{i}{}^{\dagger}(x,y)D_{F}(x-y)C^{i}(x,y)+h.c.\biggr{)}$
@@
$S=S_{KTH}+\!\int_{x}\!\biggl{(}[\overline{\psi}_{L}(x)iD\!\!\!/\!\,\,\psi_{L}(x)]+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]\biggr{)}$
@@
$\overline{g}_{0}^{2}={g_{0}^{2}}\left(1-\frac{N_{c}g_{0}^{2}}{8\pi^{2}}\right)^{-1}.$ (88)
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\overline{g}_{0}^{2}\!\int_{xy}\!\!\!\!\sqrt{N_{c}}M_{0}^{2}(H^{i\dagger}(x,y)D_{F}(x-y)[\overline{\psi}_{R}(y)\psi_{iL}(x)]_{f}+h.c.)\approx\overline{g}_{Y}\int d^{4}X\biggl{(}H^{i\dagger}(X)[\overline{\psi}_{R}(X)\psi_{iL}(X)]_{f}+h.c.\biggr{)}.$ (89)
@@
$\overline{g}_{Y}=\overline{g}_{0}^{2}\sqrt{2N_{c}/J}\;\phi(0).$ (90)
@@
$S=\!\int_{x}\!\biggl{(}[\overline{\psi}_{L}(x)iD\!\!\!/\!\,\,\psi_{L}(x)]_{f}+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]_{f}\biggr{)}$
@@
$S=\!\int_{x}\!\biggl{(}[\overline{\psi}_{L}(x)iD\!\!\!/\!\,\,\psi_{L}(x)]_{f}+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]_{f}\biggr{)}$
@@
$\sim\widetilde{Z}\partial_{X}B^{\prime i\dagger}\partial_{X}B^{\prime}_{i}-M_{0}^{2}B^{\prime i\dagger}B^{\prime}_{i}\qquad\text{where,}\qquad\widetilde{Z}=\frac{\overline{g}_{0}^{2}N_{c}}{8\pi^{2}}\ln\left(\frac{M_{0}}{m}\right),$ (94)
@@ IV.1 Physical Consequences of the Improved Yukawa Coupling
$\frac{\overline{g}_{0c}^{2}N_{c}}{8\pi^{2}}=1.06940,\qquad\text{cf. NJL model:}\;\;\frac{\overline{g}_{0c}^{2}N_{c}}{8\pi^{2}}=1.$ (96)
@@
$\phi(r)\approx\frac{ce^{-|\mu|r}}{r};\;\;\;\;\;\makebox{then we find,}\;\;\;\;\;\phi(0)=\frac{2\sqrt{2}}{\pi^{3/2}}\sqrt{\frac{|\mu|}{M_{0}}}=0.50795\sqrt{\frac{|\mu|}{M_{0}}}.$ (97)
@@
we then predict $M_{0}=6.745\;\makebox{TeV}$. [^8] This was computed in main where we quoted $5.9$ TeV, which resulted from using the NJL critical value $N_{c}g_{0c}^{2}/8\pi^{2}=1$ rather than the true $N_{c}g_{0c}^{2}/8\pi^{2}=1.0694$. However, we think these are within the uncertainty of the leading order calculations for the SKG solutions, and does not take RG running effects, e.g., logarithms, into account. Likewise, using an approximate numerical solution to the SKG equation yields $M_{0}=5.23\;\makebox{TeV}$ (this is sensitive to a fitting function, and likely underestimates $M_{0}$). Conservatively we would estimate $M_{0}$ to lie in a range of order $5<M_{0}<7$ TeV.
@@
$\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\approx\frac{1.0694}{1+1.0694}=0.51677,$ (99)
@@ IV.2 Physical Consequences of the Improved Quartic Coupling
$\lambda\approx(\overline{g}_{Y}^{4})\frac{N_{c}}{4\pi^{2}}\ln\left(\frac{M_{0}}{|\mu|}\right)\approx 0.321\;\;\;\makebox{(cf., $0.25$ experiment).}$ (101)
@@
$\lambda\approx(\overline{g}_{Y}^{4}-\overline{g}_{Y}^{2}\lambda-\lambda^{2})\frac{N_{c}}{4\pi^{2}}\ln\left(\frac{M_{0}}{|\mu|}\right)\approx 0.230\;\;\;\makebox{(cf., $0.25$ experiment).}$ (102)
@@
$\lambda\approx(\overline{g}_{Y}^{4}-\overline{g}_{Y}^{2}\lambda)\frac{N_{c}}{4\pi^{2}}\ln\left(\frac{M_{0}}{\mu}\right)\approx 0.243\;\;\;\makebox{(cf., $0.25$ experiment).}$ (103)
@@ Acknowledgments
I thank Bill Bardeen, Thom Curtright, Bogdan Dobrescu, Andreas Kronfeld and Julius Kuti for discussions and comments during the course of this work.